OpenStudy (anonymous):

Find equation of a tangent line to the graph of f(theta)=tan(theta) at point(pi/4, 1)

OpenStudy (dumbcow):

f'(x) = 1/cos^2 at x = pi/4 cos(pi/4) = sqrt2/2 f'(pi/4) = 1/(1/2) = 2 --> slope y = mx+b 1 = 2(pi/4) + b b = 1-pi/2 --> y = 2x +(1-pi/2)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yw